OpenStudy (anonymous):

Compare 4/7 and 1/6 please. c:

OpenStudy (anonymous):

Try to make the denominators (the numbers on the bottom the same) How would you do that?

OpenStudy (anonymous):

Oh. Well 7 is clearly bigger than 6, is that what you meant?

OpenStudy (anonymous):

No \[\frac{4}{7}\] and \[\frac{1}{6}\] we need to get the numbers on the bottom the same. Let's multiply each by the denominator of the other. \[\frac{4}{7} = \frac{4 \times 6}{7 \times 6} = \frac{24}{42}\]\[\frac{1}{6} = \frac{1 \times 7}{6 \times 7} = \frac{7}{42}\] Now we have the two fractions over a common divisor. So now we can simply relate 24 and 7. I know that 24 is about 3.43 times bigger than 7! Does that make sense?

OpenStudy (anonymous):

Yes So, My answer would be 4/7 > 1/6, right?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Thanks!